ServiceAgent1:

Polly and Molly are writing letters to all of the athletes they love most. They have a list of 15 athletes, and Polly has already written 2 letters. Molly has written 4 letters. Write an equation with the variable (L) and show your work to determine how many letters they have left to write.

smackz:

hehe

ServiceAgent1:

my twin brotherrr

smackz:

o;

smackz:

ily more then my self ;o

smackz:

buh lemem answa yo question

ServiceAgent1:

lol

ServiceAgent1:

ok

smackz:

15- (4+2)=L i beleive is what u can use

smackz:

wait e.e

smackz:

i might be wrong e.e

smackz:

we need experts ;c

simsharrison:

(2+4 = 6) 15-6 =??

simsharrison:

L=9 so they have 9 letters left to write.

ServiceAgent1:

so i think @simsharrison is correct ig

simsharrison:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ServiceAgent1 so i think @simsharrison is correct ig \(\color{#0cbb34}{\text{End of Quote}}\) np